Project

General

Profile

NDOS Commissioning Meeting 2011/03/21

Attendance: Jon, Chad, Kurt, Andrew, Jarek, Minerba, Susan, Mark Messier, Gary, Leon, Alena, Enrique, Peter, Athans, Sue, Rick K., Greg, Mat, Gavin, Denis, Mark Mathis

Agenda:

  1. NDOS Filling Update; open filling shifts
  2. DAQ/Electronics Update & Plans
  3. DCS Status & Plans
  4. AOB

NDOS Filling:

  • Minerba: we've been filling since last Thursday evening, went thru Saturday. Today Gavin and Mat are filling. Should complete first filling this week, possibly get to top-off. There are some open shifts: Thursday all day and Friday afternoon.

DAQ Electronics:

  • Crimping tool for noise abatement
  • APD test stand arrived at CalTech on Friday, will be installed today. 30 APDs will be used to check it out, then

DAQ Software:

  • Peter: last week there was a bit of a struggle getting (new) software working. Latest versions of DCM software. Saw increase in CPU usage (both system and user).
  • Kurt: increased buffer size to help with lost trigger messages. Seems to work on test-stand. Moving forward on this today (along with new version of DDS).
  • Progress on DCM WOFO? Sue: tried to run some tests last Thursday and Friday. Worked with Ron R on Friday. Problems with timing chain (one of DCMs was not responding to timing commands). Discovered they were using an older version of TDU firmware at test-stand. Has had a hard time reproducing the problem. Peter: Rick K. is also working on firmware re: timing. Still in progress. Peter suggests that shifters fill out a checklist for crashes.
  • Need FEB checkout procedure! Sue's validateDCM script will attempt to communicate with any attached FEBs. Will try to work on this this week.

DCS Update:

  • Chad will start wiring up RTDs this week on detector. Will need to use MAP.
  • Athans: showed

Any other business:

  • Denis: want to try to install noise abatement boards on detector this week. Will try different installation procedures, etc. Will install on "good" (well-behaved) APDs (non-decaying, etc.)
  • Mark: how often are we generating Miniboone timestamps? Denis: this weekend he processed all data from December.
  • Moving to new electronic logbook (ECL)!
  • Andrew: this is computer security week! Auditors will be coming around to check on laptops, desktops, etc. If you see people you don't know in an area where they should not be, as for identification.
  • Andrew: we are supposed to be getting USB cables to connect to console ports on DCMs sometime this week. Will need time on detector at some point.